OpenStudy (anonymous):

What is the solution set of 7x^2 + 3x = 0? {0, 3/7} {0, -3/7} {0, -4/7}

OpenStudy (anonymous):

You may factor out the "x" on left side first. Then equate each factor to zero.

OpenStudy (anonymous):

so either\[7x ^{2}=0\] or \[3x=0\]

mathslover (mathslover):

either use quadratic equation or factoring: \[\large{\color{red}{\textbf{By factoring first}}}\] \[\large{\color{blue}{7x^2+3x=0}}\] \[\large{\color{green}{x(7x+3)=0}}\] \[\large{\color{brown}{\textbf{Either x = 0 or 7x+3=0, can you find out the soln?}}}\]

OpenStudy (anonymous):

so its -3/7 and 0?